# Sard properties for polynomial maps in infinite dimension

### v2

@@ 4.5. A class of maps with the global Sard property
(244) $\sup_{x\in K}\left\lVert\left(D_{x}f\right)|_{E_{n}}-D_{\pi_{E_{n}}(x)}f_{n}\right\rVert_{\operatorname{op}}$ $\leq\sup_{x\in K}\left\lVert D_{x}f-D_{x}(f_{n}\circ\pi_{E_{n}})\right\rVert_{\operatorname{op}}$
(245) $\leq\sup_{x\in K}\sum_{k=n+1}^{\infty}\left\lVert D_{x}(p_{k}\circ\pi_{k})\right\rVert_{\operatorname{op}}$ by (215)
(246) $=\sup_{x\in K}\sum_{k=n+1}^{\infty}\left\lVert(D_{\pi_{k}(x)}p_{k})\circ\pi_{k}\right\rVert_{\operatorname{op}}$
@@
Note that, in the last line, the supremum is taken on a finite-dimensional Euclidean ball. To proceed, by Lemma 38 and (240) we have
@@
(249) $\sup_{x\in K}\left\lVert\left(D_{x}f\right)|_{E_{n}}-D_{\pi_{E_{n}}(x)}f_{n}\right\rVert_{\operatorname{op}}$ $\leq\frac{d^{2}\sqrt{m}}{r}\alpha(d,m)(1+r)^{d}\sum_{k=n+1}^{\infty}\frac{k^{d}}{q^{k}}.$
@@
(250) $\sup_{x\in K}\bigg{(}\|f(x)-f_{n}\circ\pi_{E_{n}}(x)\|+\|\left(D_{x}f\right)|_{E_{n}}-D_{\pi_{E_{n}}(x)}f_{n}\|_{\operatorname{op}}\bigg{)}\leq\left(1+\frac{d^{2}\sqrt{m}}{r}\right)\alpha(d,m)(1+r)^{d}\sum_{k=n+1}^{\infty}\frac{k^{d}}{q^{k}}.$
@@
for sufficiently large $n$, where $c\geq 1$ is a constant depending only on the fixed parameters $r,d,m$. Therefore, the main assumption (18) of Theorem D is satisfied, yielding
(252) $\dim_{e}\bigg{(}f\big{(}\operatorname{Crit}_{\nu}(f)\cap B_{H}(r)\big{)}\bigg{)}\leq\nu+\frac{\ln\beta_{0}}{\ln q},\qquad\forall\,\nu={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0},\dots,m-1.$
@@ 4.6. Sard threshold on compacts
We can now prove Theorem F, of which we recall the statement.
@@
By Theorem A this set is non-empty and $\omega_{0}(d,m)\geq\beta_{0}(d,m)^{-1}$. Hence $\omega_{0}(d,m)\in(0,1]$. Both items of Theorem 42 follow. Note that by Theorem B (more specifically, the construction of Example 34), provided that $d\geq 3$, one must have $\omega_{0}(d,m)\leq(d-1)^{-1/d}<1$, so that Item ii is non-vacuous. ∎
@@ 5. Applications to the Endpoint maps of Carnot groups
where $\mathfrak{g}_{i}\neq\{0\}$, $\mathfrak{g}_{i+1}=[\mathfrak{g}_{1},\mathfrak{g}_{i}]$ for all $i=1,\dots,s-1$ and $[\mathfrak{g}_{1},\mathfrak{g}_{s}]=\{0\}$. We set $k_{i}=\dim\mathfrak{g}_{i}$, $a_{i}=k_{1}+\dots+k_{i}$, and $k=k_{1}$, which is called the rank of the Carnot group.
@@
We recall from [Bel96, Prop. 5.18] the following result on the structure of the vector fields $X_{j}$.
@@ 5.1. Endpoint maps
Thanks to the special form of the vector fields $X_{j}$ given by Proposition 44, we can rewrite the system (266) in exponential coordinates. Adopting the notation $\gamma=(\gamma_{1},\dots,\gamma_{m})$, we have
@@
The next result (Theorem H in the Introduction) connects this framework to the previous sections.
Theorem 46.
@@
The fact that $\mathcal{E}$ is $\mathcal{C}^{1}$ (in fact smooth) and that the differential $D\mathcal{E}$ is locally Lipschitz follows from [ABB20, Prop. 8.5]. The fact that the $D\mathcal{E}$ is weakly continuous is proved in [BL17, Thm. 23]. It remains to show that for all finite dimensional linear subspaces $E$ of $H$, the restriction $\mathcal{E}|_{E}:E\to\mathbb{R}^{m}$ is a polynomial map of degree at most $s$.
@@
[…]
Recall that $Q_{ji}$ is a homogeneous polynomial of weighted degree $w_{i}-1$ and as already highlighted in the notation, depends only on the functions $\gamma_{u,\ell}(s)$ with $w_{\ell}\leq w_{i}-1=\theta$. By the induction assumption, all the functions $E\ni u\mapsto\gamma_{u,\ell}(s)$ are polynomials of degree $w_{\ell}\leq\theta$, depending continuously on $s$. By using the definition of weighted degree (see 4), it follows that the function
@@
Thanks to Theorem H we can apply the results of Section 4 to Endpoint maps of Carnot groups, obtaining Theorem I in the Introduction.
@@ 5.2. Sard property for real–analytic controls
We specialize the result in Theorem I to some concrete set of controls $K$, estimating its $n$–width.
@@
(273) $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r):=\bigg{\{}u\in L^{2}(I,\mathbb{R}^{k})\,\bigg{|}\,\textrm{$u$ is real--analytic with radius of convergence strictly greater than }r\bigg{\}},$
@@
Theorem 47.
Fix $I=[0,1]$, $\ell\in\mathbb{N}$ and $r>1$. Let $V=\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$. Then the set $K_{V}:=\{u\in V\mid\|u\|_{V}\leq 1\}$ is compact in $L^{2}(I,\mathbb{R}^{k})$, and for its $n$–width it holds:
(278) $\Omega_{n}(K_{V},L^{2}(I,\mathbb{R}^{k}))\leq\frac{(k\ell)^{1/2}}{\ln r}\left(\frac{1}{r}\right)^{\left\lfloor\frac{n}{k\ell}\right\rfloor}.$
@@
We start by proving the case $\ell=1$. Fix $u=(u_{1},\dots,u_{k})\in K_{V}$. Since $r>1$, any $u_{j}$ for $j=1,\dots,k$ has an expansion in power series on the complex unit ball centered at the origin, denoted by $B_{\mathbb{C}}(1)$. In particular, we have for any $t\in I$
@@
[…]
It follows that for all $u\in K_{V}$ it holds
@@
(286) $\Omega_{kN}(K_{V},L^{2}(I,\mathbb{R}^{k}))\leq\frac{k^{1/2}}{\ln r}\left(\frac{1}{r}\right)^{N}.$
For $n\in\mathbb{N}$ we consider $N=\left\lfloor\frac{n}{k}\right\rfloor$. By Definition 26, the $n$–width is non-increasing as a function of $n\in\mathbb{N}$, hence by (286) we have
(287) $\Omega_{n}(K_{V},L^{2}(I,\mathbb{R}^{k}))\leq\Omega_{kN}(K_{V},L^{2}(I,\mathbb{R}^{k}))\leq\frac{k^{1/2}}{\ln r}\left(\frac{1}{r}\right)^{\left\lfloor\frac{n}{k}\right\rfloor}.$
This proves the estimate on the $n$–width per $\ell=1$.
@@
where $I_{a}$ are the intervals of the decomposition (275), and $\mathbbm{1}_{I_{a}}$ are the corresponding characteristic functions. By definition, each $u|_{I_{a}}$ is the restriction to $I_{a}$ of a real–analytic function with radius of convergence $r_{a}\geq r>1$. We can expand each $u|_{I_{a}}$ in Taylor series centered at the lower bound of $I_{a}$. Then, in order to obtain the finite-dimensional approximation, we truncate the series as in the previous case, repeating analogous estimates for the remainder (taking into account the length of the intervals $|I_{a}|=1/\ell$). This concludes the estimate on the $n$–width for general $\ell$.
To prove the compactness assertion, observe that there is a linear immersion (given by the inclusion) $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)\hookrightarrow L^{2}(I,\mathbb{R}^{k})$ and it holds $\|u\|_{L^{2}(I,\mathbb{R}^{k})}\leq\|u\|_{C^{\omega}(I,\mathbb{R}^{k};r,\ell)}$ for all $u\in C^{\omega}(I,\mathbb{R}^{k};r,\ell)$. In particular $K_{V}$ is bounded in $L^{2}(I,\mathbb{R}^{k})$ and since its $n$–width tends to zero as $n\to\infty$ it is also compact by Proposition 27. ∎
We can now prove Theorem J in the Introduction, of which we recall the statement.
@@
Let $\mathbb{G}$ be a Carnot group of topological dimension $m$, step $s$, and rank $k$. Given $\ell\in\mathbb{N}$, there exists $r=r(m,s,k,\ell)>0$ such that, letting $V=\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$, with $I=[0,1]$, it holds
(289) $\mu\big{(}\mathcal{E}(\mathrm{Crit}(\mathcal{E})\cap V)\big{)}=\mu\big{(}\mathcal{E}(\mathrm{Crit}(\mathcal{E}|_{V})\big{)}=0.$
Namely, the Sard property holds on the space of piecewise real–analytic controls with radius of convergence $>r$ and with $\ell$ pieces. Furthermore, letting $W=\mathcal{C}^{\omega}(I,\mathbb{R}^{k};\infty,\ell)$, then it holds
(290) $\dim_{\mathcal{H}}\big{(}\mathcal{E}(\mathrm{Crit}(\mathcal{E})\cap W)\big{)}\leq\dim_{\mathcal{H}}\big{(}\mathcal{E}(\mathrm{Crit}(\mathcal{E}|_{W})\big{)}\leq m-1.$
@@
Observe that $\mathcal{E}\in\mathscr{P}^{m}_{s}(L^{2}(I,\mathbb{R}^{k}))$ by Theorem H. Let $V=\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$, and $K_{V}$ be its unit ball with respect to the norm (277). By Theorem 47, for any $r>1$ we have
(291) $\limsup_{n\to\infty}\Omega_{n}(K_{V},L^{2}(I,\mathbb{R}^{k}))^{1/n}\leq r^{-\tfrac{1}{k\ell}}.$
We can apply Theorem I, observing that $\mathrm{span}(K_{V})=V$. ∎
@@ 5.3. Surjectivity of the Endpoint map on finite-dimensional spaces of controls
The Endpoint maps of sub-Riemannian manifolds are surjective when restricted to piecewise constant controls: this follows from the proof of the Rashevskii-Chow theorem in [ABB20]. In this section, we prove that in Carnot groups the Endpoint maps are surjective also when they are restricted to the space of controls which consists in the set of polynomial maps of some large enough fixed degree (which depends on the Carnot group). The proof is obtained by using a quantitative version of the inverse function theorem, as it can be found in [Cla76]. For completeness we present and prove in our setting the statements we need.
@@
From [Cla76, Lemma 3] we obtain the following technical lemma.
@@
is convex. We now prove that its distance from 00 is at least $\frac{\sigma(M_{0})}{2}$. Indeed, by (293), for any $M$ with $\|M-M_{0}\|_{\operatorname{op}}\leq\tfrac{\sigma(M_{0})}{2}$ we obtain $\sigma(M)>\frac{\sigma(M_{0})}{2}$. Therefore,
@@
From [Cla76, Lemma 4] we obtain the following “quantitative injectivity” lemma.
@@
[…]
We define $v_{x,y}:=\frac{x-y}{\|x-y\|}\in\mathbb{R}^{m}$, and we apply Lemma 49 to $v=v_{x,y}$ and $M_{0}=D_{x_{0}}f$, which is invertible by hypothesis. We denote by $w_{x,y}\in\mathbb{R}^{m}$ the provided unit vector, such that
@@
[…]
@@
We conclude with the following inclusion, analogous to [Cla76, Lemma 5], providing the “quantitative surjectivity” counterpart of the previous statement. We report the proof for completeness.
@@
In the same setting of Lemma 50, the following inclusion holds
@@
Set $r=r_{f,x_{0}}$ and $\sigma=\sigma_{f,x_{0}}$. Let $y\in B_{\mathbb{R}^{m}}\left(g(x_{0}),\frac{r\sigma}{8}\right)$. Let $x$ be a minimum of $\|y-g(\cdot)\|^{2}$ on $B_{\mathbb{R}^{n}}(x_{0},r)$. We claim that $x\in\mathrm{int}B_{\mathbb{R}^{n}}(x_{0},r)$. Otherwise, using Lemma 50, we have
@@
which is a contradiction. Thus $x$ yields a local minimum for the function $\|y-g(\cdot)\|^{2}$, and consequently $\nabla_{x}\|y-g(\cdot)\|^{2}=-2D_{x}g\cdot(y-g(x))=0$. Note that by (301) and the fact that the singular values are $1$-Lipschitz (see Lemma 17) it follows that $\sigma(D_{x}g)>\tfrac{\sigma}{2}$, hence $D_{x}g$ is invertible. Hence $y=g(x)$. ∎
We can now state and prove the surjectivity property of the Endpoint maps of Carnot groups.
Theorem 52.
@@
(311) ($\mathcal{E}\left(\{u\in L^{2}(I,\mathbb{R}^{k})\mid u\text{ is a polynomial map with components of degrees at most }d_{\mathbb{G}}\}\right)=\mathbb{G}.$
@@
The classical proof of the Rashevskii-Chow theorem provides $u_{0}\in L^{2}(I,\mathbb{R}^{k})$ such that $D_{u_{0}}\mathcal{E}$ is surjective and $\mathcal{E}(u_{0})=0$ (see e.g. [ABB20, Sec. 3.2]). In particular there exist $w_{1},\dots,w_{m}\in\ker(D_{u_{0}}\mathcal{E})^{\perp}$ such that the vectors $D_{u_{0}}\mathcal{E}(w_{1}),\dots,D_{u_{0}}\mathcal{E}(w_{m})$ are linearly independent. Now we consider the map $f:\mathbb{R}^{m}\to\mathbb{G}\simeq\mathbb{R}^{m}$ defined as
@@
where $q_{0},p_{1},\dots,p_{m}\in L^{2}(I,\mathbb{R}^{k})$ are suitably chosen polynomial maps, in such a way that the image of $g$ stills contains a neighbourhood of 00. This will be done by applying Lemma 51, as we now explain.
@@
Now we prove that we can find polynomial maps $p_{1},\dots,p_{m}$ and a neighbourhood $W\subset\mathbb{R}^{m}$ of 00 such that, setting $\sigma_{f,0}=\sigma(D_{x_{0}}f)$, it holds

### v3

@@ 4.5. A class of maps with the global Sard property
(247) $\sup_{x\in B}\left\lVert\left(D_{x}f\right)|_{E_{n}}-D_{\pi_{E_{n}}(x)}f_{n}\right\rVert_{\operatorname{op}}$ $\leq\sup_{x\in B}\left\lVert D_{x}f-D_{x}(f_{n}\circ\pi_{E_{n}})\right\rVert_{\operatorname{op}}$
(248) $\leq\sup_{x\in B}\sum_{k=n+1}^{\infty}\left\lVert D_{x}(p_{k}\circ\pi_{k})\right\rVert_{\operatorname{op}}$ by (219)
(249) $=\sup_{x\in B}\sum_{k=n+1}^{\infty}\left\lVert(D_{\pi_{k}(x)}p_{k})\circ\pi_{k}\right\rVert_{\operatorname{op}}$
@@
Note that, in the last line, the supremum is taken on a finite-dimensional Euclidean ball. To proceed, by Lemma˜43 and (243) we have
@@
(252) $\sup_{x\in B}\left\lVert\left(D_{x}f\right)|_{E_{n}}-D_{\pi_{E_{n}}(x)}f_{n}\right\rVert_{\operatorname{op}}$ $\leq\frac{d^{2}\sqrt{m}}{r}\alpha(d,m)(1+r)^{d}\sum_{k=n+1}^{\infty}\frac{k^{d}}{q^{k}}.$
@@
(253) $\sup_{x\in B}\bigg(\|f(x)-f_{n}\circ\pi_{E_{n}}(x)\|+\|\left(D_{x}f\right)|_{E_{n}}-D_{\pi_{E_{n}}(x)}f_{n}\|_{\operatorname{op}}\bigg)\leq\left(1+\frac{d^{2}\sqrt{m}}{r}\right)\alpha(d,m)(1+r)^{d}\sum_{k=n+1}^{\infty}\frac{k^{d}}{q^{k}}.$
@@
for sufficiently large $n$, where $c\geq 1$ is a constant depending only on the fixed parameters $r,d,m$. Therefore, the main assumption (19) of Theorem˜D is satisfied, yielding
(255) $\dim_{e}\bigg(f\big(\operatorname{Crit}_{\nu}(f)\cap B_{H}(r)\big)\bigg)\leq\nu+\frac{\ln\beta_{0}}{\ln q},\qquad\forall\,\nu=0,\dots,m-1.$
@@ 4.6. Sard threshold on compacts
We can now prove Theorem˜F, of which we recall the statement.
@@
By Theorem˜A this set is non-empty and $\omega_{0}(d,m)\geq\beta_{0}(d,m)^{-1}$. Hence $\omega_{0}(d,m)\in(0,1]$. Both items of Theorem˜47 follow. Note that by Theorem˜B (more specifically, the construction of Example˜38), provided that $d\geq 3$, one must have $\omega_{0}(d,m)\leq(d-1)^{-1/d}<1$, so that Item˜ii is non-vacuous. ∎
@@ 5. Applications to the Endpoint maps of Carnot groups
where $\mathfrak{g}_{i}\neq\{0\}$, $\mathfrak{g}_{i+1}=[\mathfrak{g}_{1},\mathfrak{g}_{i}]$ for all $i=1,\dots,s-1$ and $\mathfrak{g}_{1},\mathfrak{g}_{s}]=\{0\}$. We set $k_{i}=\dim\mathfrak{g}_{i}$, $a_{i}=k_{1}+\dots+k_{i}$, and $k=k_{1}$, which is called the rank of the Carnot group.
@@
We recall from [Bellaiche, Prop. 5.18] the following result on the structure of the vector fields $X_{j}$.
@@ 5.1. Endpoint maps
Thanks to the special form of the vector fields $X_{j}$ given by Proposition˜50, we can rewrite the system (269) in exponential coordinates. Adopting the notation $\gamma=(\gamma_{1},\dots,\gamma_{m})$, we have
@@
The next result (Proposition˜H in the Introduction) connects this framework to the previous sections.
Proposition 52.
@@
The fact that $\mathcal{E}$ is $\mathcal{C}^{1}$ (in fact smooth) and that the differential $D\mathcal{E}$ is locally Lipschitz follows from [ABB, Prop. 8.5]. The fact that the $D\mathcal{E}$ is weakly continuous is proved in [BoarLer17, Thm. 23]. It remains to show that for all finite dimensional linear subspaces $E$ of $H$, the restriction $\mathcal{E}|_{E}:E\to\mathbb{R}^{m}$ is a polynomial map of degree at most $s$.
@@
(273) $\gamma_{u,i}(t)=\sum_{j=1}^{k}\int_{0}^{t}u_{j}(s)Q_{ji}(\gamma_{u,1}(s),\dots,\gamma_{u,k_{w_{i}-1}}(s))\,ds.$
Recall that $Q_{ji}$ is a homogeneous polynomial of weighted degree $w_{i}-1$ and as already highlighted in the notation, depends only on the functions $\gamma_{u,\ell}(s)$ with $w_{\ell}\leq w_{i}-1=\theta$. By the induction assumption, all the functions $E\ni u\mapsto\gamma_{u,\ell}(s)$ are polynomials of degree $w_{\ell}\leq\theta$, depending continuously on $s$. By using the definition of weighted degree (see Remark˜49), it follows that the function
@@
Thanks to Proposition˜H we can apply the results of Section˜4 to Endpoint maps of Carnot groups, obtaining Corollary˜I in the Introduction.
@@ 5.2. Sard property for real–analytic controls
We specialize the result in Corollary˜I to some concrete set of controls $K$, estimating its $n$-width.
@@
[…]
@@
Proposition 53.
Fix $I=[0,1]$, $\ell\in\mathbb{N}$ and $r>1$. Let $Y=\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$. Then the set $B_{Y}:=\{u\in V\mid\|u\|_{Y}\leq 1\}$ is precompact in $L^{2}(I,\mathbb{R}^{k})$, and for its $n$-width it holds:
(281) $\Omega_{n}(B_{Y},L^{2}(I,\mathbb{R}^{k}))\leq\frac{(k\ell)^{1/2}}{\ln r}\left(\frac{1}{r}\right)^{\left\lfloor\frac{n}{k\ell}\right\rfloor}.$
@@
We start by proving the case $\ell=1$. Fix $u=(u_{1},\dots,u_{k})\in B_{Y}$. Since $r>1$, any $u_{j}$ for $j=1,\dots,k$ has an expansion in power series on the complex unit ball centered at the origin, denoted by $B_{\mathbb{C}}(1)$. In particular, we have for any $t\in I$
@@
(286) $\|u_{j}-u_{j,N}\|_{L^{2}(I)}\leq\left\lVert u_{j}-u_{j,N}\right\rVert_{L^{\infty}(I)}\leq\sum_{h=N}^{\infty}\frac{|u_{j}^{(h)}(0)|}{h!}\leq\sum_{h=N}^{\infty}\frac{1}{r^{h}}\leq\int_{N}^{\infty}\frac{1}{r^{x}}\,dx=\frac{r^{-N}}{\ln r}.$
It follows that for all $u\in B_{Y}$ it holds
@@
(289) $\Omega_{kN}(B_{Y},L^{2}(I,\mathbb{R}^{k}))\leq\frac{k^{1/2}}{\ln r}\left(\frac{1}{r}\right)^{N}.$
For $n\in\mathbb{N}$ we consider $N=\left\lfloor\frac{n}{k}\right\rfloor$. By Definition˜30, the $n$-width is non-increasing as a function of $n\in\mathbb{N}$, hence by (289) we have
(290) $\Omega_{n}(B_{Y},L^{2}(I,\mathbb{R}^{k}))\leq\Omega_{kN}(B_{Y},L^{2}(I,\mathbb{R}^{k}))\leq\frac{k^{1/2}}{\ln r}\left(\frac{1}{r}\right)^{\left\lfloor\frac{n}{k}\right\rfloor}.$
This proves the estimate on the $n$-width for $\ell=1$.
@@
where $I_{a}$ are the intervals of the decomposition (278), and $\mathbbm{1}_{I_{a}}$ are the corresponding characteristic functions. By definition, each $u|_{I_{a}}$ is the restriction to $I_{a}$ of a real–analytic function with radius of convergence $r_{a}\geq r>1$. We can expand each $u|_{I_{a}}$ in Taylor series centered at the lower bound of $I_{a}$. Then, in order to obtain the finite-dimensional approximation, we truncate the series as in the previous case, repeating analogous estimates for the remainder (taking into account the length of the intervals $|I_{a}|=1/\ell$). This concludes the estimate on the $n$-width for general $\ell$.
To prove the compactness assertion, observe that there is a linear immersion (given by the inclusion) $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)\hookrightarrow L^{2}(I,\mathbb{R}^{k})$ and it holds $\|u\|_{L^{2}(I,\mathbb{R}^{k})}\leq\|u\|_{C^{\omega}(I,\mathbb{R}^{k};r,\ell)}$ for all $u\in C^{\omega}(I,\mathbb{R}^{k};r,\ell)$. In particular $B_{Y}$ is bounded in $L^{2}(I,\mathbb{R}^{k})$ and since its $n$-width tends to zero as $n\to\infty$ it is also precompact by Proposition˜31. ∎
We can now prove Theorem˜J in the Introduction, of which we recall the statement.
@@
Let $\mathbb{G}$ be a Carnot group of topological dimension $m$, step $s$, and rank $k$. Let $I=[0,1]$. For $\ell\in\mathbb{N}$, there exists $r=r(m,s,k,\ell)>1$ such that
(292) $\mu\big(\mathcal{E}(\mathrm{Crit}(\mathcal{E})\cap\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell))\big)=\mu\big(\mathcal{E}(\mathrm{Crit}(\mathcal{E}|_{\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)})\big)=0.$
Namely, the Sard property holds on the space of piecewise real–analytic controls with radius of convergence >r>r and with $\ell$ pieces. Furthermore it holds
(293) $\dim_{\mathcal{H}}\big(\mathcal{E}(\mathrm{Crit}(\mathcal{E})\cap\mathcal{C}^{\omega}(I,\mathbb{R}^{k};\infty,\ell))\big)\leq\dim_{\mathcal{H}}\big(\mathcal{E}(\mathrm{Crit}(\mathcal{E}|_{\mathcal{C}^{\omega}(I,\mathbb{R}^{k};\infty,\ell)})\big)\leq m-1.$
@@
Observe that $\mathcal{E}\in\mathscr{P}^{m}_{s}(L^{2}(I,\mathbb{R}^{k}))$ by Proposition˜H. Let $Y:=\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$. Let $B_{Y}$ be the unit ball of $Y$ with respect to the norm (280), namely $B_{Y}=\{u\in Y\mid\|u\|_{Y}\leq 1\}$. By Propositions˜53 and 31, for any $r>1$ we have that $K:=\overline{B_{Y}}$ is compact (we stress that the closure is taken in the $L^{2}(I,\mathbb{R}^{k})$ topology, and not w.r.t. the norm $\|\cdot\|_{Y}$), and it holds
(294) $\limsup_{n\to\infty}\Omega_{n}(K,L^{2}(I,\mathbb{R}^{k}))^{1/n}=\limsup_{n\to\infty}\Omega_{n}(B_{Y},L^{2}(I,\mathbb{R}^{k}))^{1/n}\leq r^{-\tfrac{1}{k\ell}},$
where we used Proposition˜32 for the first equality. We can apply Item˜(i) of Corollary˜I: for $r>r(m,s,k,\ell):=\beta_{0}(s,m)^{k\ell}$ it holds
(295) $\mu\big(\mathcal{E}(\mathrm{Crit}(\mathcal{E})\cap\mathrm{span}(K))\big)=\mu\big(\mathcal{E}(\mathrm{Crit}(\mathcal{E}|_{\mathrm{span}(K)})\big)=0.$
By Lemma˜41, the above equalities remain valid replacing $X:=\mathrm{span}(K)$ with any linear subspace of the latter, dense in it w.r.t. the $L^{2}(I,\mathbb{R}^{k})$ topology. In particular, this is the case for $Y$, observing that $Y=\mathrm{span}(B_{Y})$ and thus $\bar{Y}=\overline{\mathrm{span}(B_{Y})}\supseteq\mathrm{span}(\overline{B_{Y}})=\mathrm{span}(K)$.
The same argument via Item˜(ii) of Corollary˜I yields the statement for the space $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};\infty,\ell)$. ∎
@@ 5.3. Surjectivity of the Endpoint map on finite-dimensional spaces of controls
The Endpoint maps of sub-Riemannian manifolds are surjective when restricted to piecewise constant controls: this follows from the proof of the Rashevskii-Chow theorem in [ABB]. In this section, we prove that in Carnot groups the Endpoint maps are surjective also when they are restricted to the space of controls which consists in the set of polynomial maps of some large enough fixed degree (which depends on the Carnot group). The proof is obtained by using a quantitative version of the inverse function theorem, as it can be found in [ClarkeInvFunc]. For completeness we present and prove in our setting the statements we need.
@@
From [ClarkeInvFunc, Lemma 3] we obtain the following technical lemma.
@@
is convex. We now prove that its distance from 0 is at least $\frac{\sigma(M_{0})}{2}$. Indeed, by (297), for any $M$ with $\|M-M_{0}\|_{\operatorname{op}}\leq\tfrac{\sigma(M_{0})}{2}$ we obtain $\sigma(M)>\frac{\sigma(M_{0})}{2}$. Therefore,
@@
From [ClarkeInvFunc, Lemma 4] we obtain the following “quantitative injectivity” lemma.
@@
(306) $g(x)-g(y)=\|x-y\|\int_{0}^{1}D_{(1-t)y+tx}g\cdot\frac{x-y}{\|x-y\|}\,dt.$
We define $v_{x,y}:=\frac{x-y}{\|x-y\|}\in\mathbb{R}^{m}$, and we apply Lemma˜55 to $v=v_{x,y}$ and $M_{0}=D_{x_{0}}f$, which is invertible by hypothesis. We denote by $w_{x,y}\in\mathbb{R}^{m}$ the provided unit vector, such that
@@
(308) $\|g(x)-g(y)\|\geq\|x-y\|\left\langle w_{x,y},\int_{0}^{1}D_{(1-t)y+tx}g\cdot\frac{x-y}{\|x-y\|}\,dt\right\rangle\geq\frac{\sigma_{f,x_{0}}}{2}\|x-y\|,$
@@
We conclude with the following inclusion, analogous to [ClarkeInvFunc, Lemma 5], providing the “quantitative surjectivity” counterpart of the previous statement. We report the proof for completeness.
@@
In the same setting of Lemma˜56, the following inclusion holds
@@
Set $r=r_{f,x_{0}}$ and $\sigma=\sigma_{f,x_{0}}$. Let $y\in B_{\mathbb{R}^{m}}\left(g(x_{0}),\frac{r\sigma}{8}\right)$. Let $x$ be a minimum of $\|y-g(\cdot)\|^{2}$ on $B_{\mathbb{R}^{n}}(x_{0},r)$. We claim that $x\in\mathrm{int}B_{\mathbb{R}^{n}}(x_{0},r)$. Otherwise, using Lemma˜56, we have
@@
which is a contradiction. Thus $x$ yields a local minimum for the function $\|y-g(\cdot)\|^{2}$, and consequently $\nabla_{x}\|y-g(\cdot)\|^{2}=-2D_{x}g\cdot(y-g(x))=0$. Note that by (305) and the fact that the singular values are $1$-Lipschitz (see Lemma˜18) it follows that $\sigma(D_{x}g)>\tfrac{\sigma}{2}$, hence $D_{x}g$ is invertible. Hence $y=g(x)$. ∎
We can now state the surjectivity property of the Endpoint maps of Carnot groups.
Proposition 58.
@@
[…]
@@
The classical proof of the Rashevskii-Chow theorem provides $u_{0}\in L^{2}(I,\mathbb{R}^{k})$ such that $D_{u_{0}}\mathcal{E}$ is surjective and $\mathcal{E}(u_{0})=0$ (see e.g. [ABB, Sec. 3.2]). In particular there exist $w_{1},\dots,w_{m}\in\ker(D_{u_{0}}\mathcal{E})^{\perp}$ such that the vectors $D_{u_{0}}\mathcal{E}(w_{1}),\dots,D_{u_{0}}\mathcal{E}(w_{m})$ are linearly independent. Now we consider the map $f:\mathbb{R}^{m}\to\mathbb{G}\simeq\mathbb{R}^{m}$ defined as
@@
where $q_{0},p_{1},\dots,p_{m}\in L^{2}(I,\mathbb{R}^{k})$ are suitably chosen polynomial maps, in such a way that the image of $g$ stills contains a neighbourhood of 0. This will be done by applying Lemma˜57, as we now explain.
@@
Now we prove that we can find polynomial maps $p_{1},\dots,p_{m}$ and a neighbourhood $W\subset\mathbb{R}^{m}$ of 0 such that, setting $\sigma_{f,0}=\sigma(D_{x_{0}}f)$, it holds